Let x ,y be the roots of the equation ax^2+bx+c=0 and x1 and -y be the roots of the equation a1x^2+b1x+c1=0,then find a quadratic equation with roots x and x1

Grade:10

1 Answers

Ritz
6 years ago
If a and b are the 0s of a quadratic polynomial f(x). Then ,the polynomial f(x) is given by ,F(x)=k(x^2-,(a+b)x+abGiven a=x and b=x1Since,a+b=x+x1ab=xx1Therefore, k[x^2-(x+x1)x+xx1]=K(x^2-x^2-xx1+xx1)K(0)=0Hence,we get that no any quadratic eq will be form through these roots